OpenStudy (anonymous):

Express the complex number in trigonometric form. -4

OpenStudy (anonymous):

4(cos 270° + i sin 270°) 4(cos 90° + i sin 90°) 4(cos 0° + i sin 0°) 4(cos 180° + i sin 180°)

zepdrix (zepdrix):

\[\Large\rm z=-4+0\mathcal i\]This complex number can be written in the form:\[\Large\rm z=r(\cos \theta+\mathcal i \sin \theta)\]Where,\[\Large\rm r=\sqrt{(-4)^2+0^2}\]and,\[\Large\rm \theta=\arctan\left(\frac{0}{-4}\right)\]We don't need to get that fancy for this problem though. Simply graphing the point should be enough to see what's going on.

zepdrix (zepdrix):